Telescope and AGw Commissioning at LFBG with IRTC - 27-28 April 2009

Observers: D.Thompson, J. Heidt (Tucson)
Telescope Operator: D.Gonzalez-Huerta
Software Support: (phone) Telescope Support:

Summary

  • Lost most of night to heavy clouds
  • New dice5 script in IRTC package for repeated offset tests
  • Timed reads/writes from/to /Repository from a local scratch disk on lbtmu05
    • 2365 MB of data transferred (IRTC datacubes, 10 files, 16 to 313 MB)
    • Iterated 5 times each direction: date ; cp ; date
    • READS (copy from NFS to the scratch disk) averaged 20.67 MB/s
    • WRITES (copy from scratch to NFS) averaged 10.38 MB/s
  • A TRACK mode preset hung at/after "wait for elevation complete" at 03:51:48.382, OSS?
  • Dropped into "none/cancelled" mode spontaneously several times.

Details

Sunset: cloudy on summit.

23:00 Opened dome, it has cleared off sufficiently to work. Tout=3.2, M1=0.9 RH=30.8% wind gusting to 12m/s.

23:31 Seeing is very bad (2 arcsec) and M5_1311 is not even close to saturating on 2s acquisition images. We will run wfsing here but it is still too cloudy to do anything. IE=-91, CA=43.

23:42 pointto M5_0547 to check if we got the right star the first time. No, or the clouds are too thick to see anything. Tried Arcturus...nothing. Resetting encoders and pointing model before trying again. David reports only ~4 stars visible.

01:30 Encoder reset completed with difficulty. Another ptspiral on Arcturus showed nothing. David reports "no stars visible" outside. Closing dome at 01:56 MST.

Developed and tested (as far as possible) a "dice5" script in the IRTC package for doing repeated offsetting tests with some flexibility. You can give variable offsets in X and Y, use RADEC or DETXY offsets, set the takeseq exposure time and number of exposures, and set the number of iterations through the dice5 pattern. I tested it closed-dome with 30x2sec sequence images (to try to give the WFS time on sky to cycle each round, may need 45x2sec) and it took 6 minutes per cycle through the dice5 pattern.

Tested read/write speed from the nfs disk (/Repository) to a local scratch disk on lbtmu05. The data copied were 10 fits datacubes from last night totalling 2365MB (one 16MB file, three 157MB files, and six 313MB files). A simple "date ; cp ; date" command line was used and the elapsed time used to calculate a transfer rate. I iterated five times on nfs->scratch and scratch->nfs.

iteration read from nfs write to nfs
1 25.93 MB/s 10.03 MB/s
2 17.68 MB/s 10.06 MB/s
3 21.30 MB/s 10.46 MB/s
4 19.04 MB/s 10.65 MB/s
5 19.42 MB/s 10.71 MB/s
avg 20.67 MB/s 10.38 MB/s

03:42 There are "a few stars" visible, so we are opening up again. IR satellite looks at least mostly clear at this point.

03:50:23.308 pointto Vega TRACK mode at PA=0
03:51:48.382 2009 Preset command wait for elevation complete for left IRTC
We appear to be "on source" and tracking, but the preset is not completing in a timely fashion.
I had to ^C out of the iraf script after aborting the preset (after 6+ minutes). A subsequent offset worked.
03:59:00.502 preset again in TRACK mode, but it hung in the same place. Tried restarting GCS at 04:01:23 (maybe problems clearing AGw stages on TRACK mode preset?)...nope. Tried homing the motors on AGw...nope. Stop and restart IIF.

04:07:14.472 resend preset...same problem. Stop and restart PSF (although log messages look OK, collimation was executed after the wait for elevation complete). This did apparently cause my preset in iraf to fail.
Tue Apr 28 04:11:19.101 2009  Preset alert for left IRTC mode TRACK: OK
Tue Apr 28 04:11:19.155 2009  Preset alert for left IRTC mode TRACK: Invalid buffer: RPCERROR
Tue Apr 28 04:11:19.207 2009  Preset command failed for left IRTC mode TRACK from IIF

04:13:10.411 preset sent again...nope. Same problem. In the PSF syslog entries, it appears to be doing things as expected:
Apr 28 04:13:24 lbtmu101 LBT_PSFL: PSFL:PrimaryMirror:adjustMirrorCollimation executed
Apr 28 04:13:24 lbtmu101 LBT_PSFL: PSFL:PSFSecondaryMirror:adjustMirrorCollimation executed
Apr 28 04:13:24 lbtmu101 LBT_PSFL: PSFL:PSFTertiaryMirror:adjustMirrorCollimation executed
Apr 28 04:13:24 lbtmu101 LBT_PSFL: Setting up block for handle 1: block:1:0
Apr 28 04:13:24 lbtmu101 LBT_PSFL: waiting for handle block:1:0 0 1
Apr 28 04:13:24 lbtmu101 LBT_PSFL: ACTOPT: Primary collimated
Apr 28 04:13:24 lbtmu101 LBT_PSFL: ACTOPT: Secondary collimated
Apr 28 04:13:26 lbtmu101 LBT_PSFL: PSFL:PSFSecondaryMirror:adjustMirrorCollimation execution time: 2541 ms
Apr 28 04:13:29 lbtmu101 LBT_PSFL: PSFL:PrimaryMirror:adjustMirrorCollimation execution time: 5255 ms
Apr 28 04:13:29 lbtmu101 LBT_PSFL: wait done for handle block:1:0 0 1
Apr 28 04:13:29 lbtmu101 LBT_PSFL: Setting up block for handle 3: block:2:0
Apr 28 04:13:29 lbtmu101 LBT_PSFL: waiting for handle block:2:0 0 3
but M3 does not appear to the completing its collimation request from PSF. Stop and restart OSS and:
Apr 28 04:16:28 lbtmu101 LBT_PSFL: PSFL:PSFTertiaryMirror:adjustMirrorCollimation execution time: 184141 ms
Apr 28 04:16:28 lbtmu101 LBT_PSFL: wait done for handle block:2:0 0 3
Apr 28 04:16:28 lbtmu101 LBT_networkserver: dropping OSS
at which point the iraf preset failed and the IIFGUI noted:
 Tue Apr 28 04:16:28.337 2009  TCS authorize alert for left IRTC at bentGregorianFront focus: OK
Tue Apr 28 04:16:28.390 2009  TCS authorize alert for left IRTC at bentGregorianFront focus: Invalid buffer: RPCERROR
Tue Apr 28 04:16:28.422 2009  Preset alert for left IRTC mode TRACK: OK
Tue Apr 28 04:16:28.442 2009  TCS authorize command failed for left IRTC at bentGregorianFront focus
Tue Apr 28 04:16:28.476 2009  Preset alert for left IRTC mode TRACK: Invalid buffer: RPCERROR
Tue Apr 28 04:16:28.528 2009  Preset command failed for left IRTC mode TRACK from IIF

04:18:40.741 preset sent again...worked. OSS must have been hung at some place, but this was not obvious. 30 minutes lost.

We found a bright star with ptspiral at IE=-180, CA=-39, but it was moving fairly quickly. We were at 84 degrees EL and at transit. Moving to a source at lower EL.

04:32:33.265 pointto M5_0707 (18:51 +59). EL~63 degrees. Not found with ptreset tyen ptspiral.

04:50:04.199 Trying Jupiter (EL~28 and rising). D=37 arcsec. Ptreset and full ptspiral with IRTC in movie mode, 0.5 sec exposures. Nothing. There is something really wrong with the telescope pointing (encoders??). See critical Issue Trak #2068 Lost 1 hour to this.

-- DavidThompson - 27 Apr 2009
Topic revision: r5 - 28 Apr 2009, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback